# Comprehensive Evaluation of Prototype Neural Networks

### v2

@@ 4.1 Prototype Visualization
The post-hoc visualization method is not bound to the individual models, allowing us to select the most suitable method for the networks. The original implementation of ProtoPNet employs a simple model-agnostic upscaling method, which was also used by ProtoPool and PIPNet. This method upscales the similarity map of a prototype to the size of the input using a cubic interpolation to visualize the similarity pattern of the prototype within the image. A key disadvantage is that it relies solely on the similarity map and assumes the preservation of spatial information in the feature extraction process. There are other model-specific methods that can better visualize the feature extraction process of the model. Based on these arguments and the findings from Xu-Darme et al.[61], which compared various visualization methods, we selected the Prototypical Relevance Propagation (PRP) method from Gautam et al.[62] for visualizing the prototypes (see Figure 1).
@@ 5 Evaluation Metrics
In this section, we discuss the techniques and metrics used to evaluate the part-based prototype models. Following the Co-12 properties from Nauta et al. [19] as a guideline, we choose metrics to evaluate the output-completeness, continuity, contrastivity, covariate complexity, and compactness.
@@
We use a total of 22 metrics, including 13 metrics that are novel to this domain, an overview of the metrics categorized by the evaluation techniques is shown in Table 1. The coverage of the model components by the metrics is shown in Figure 1, with additional information if the metrics focus on spatial, activation, or other features.
@@ 5.1 Output-Completeness Evaluation
The output-completeness evaluation focuses on the visualization method’s ability to highlight all relevant parts in the input image. To evaluate the output completeness of the PRP method, we consider the study by Sacha et al. [63], which evaluates the spatial misalignment of visualization methods. The introduced metrics measure the location change in the visualization VLC and the prototype activation change PSC when a perturbation is applied to the input image. We extend this list by including the activation change of the visualization VAC, the location change of the maximum activation in the feature map PLC, as well as the general location and activation change of the similarity pattern, PALC and PAC, respectively, thereby achieving comprehensive coverage of the model components (see Figure 1). Some metrics from Sacha et al. [63] have been renamed for consistency in terminology.
Consider an image $x$ and the perturbed counterpart $\overline{x}_{i}$, created based on the visualization of the prototype $p_{i}\in P$. Let $v_{p_{i}}(x)$ be the saliency map of prototype $p_{i}$. Each saliency map is pre-processed and only contains pixels with a relevance value above the 95th percentile. Let $b_{p_{i}}(x)$ be the bounding box around this activated region of the saliency map. We perturb an image $x$ to $\overline{x}_{i}$ by adding a Gaussian noise mask with a standard deviation of $0.05$ to the pixels around the bounding box; see Figure 5 for a process visualization. Further, let $a_{p_{i}}(x)$ be the similarity map and $s_{p_{i}}(x)$ the similarity score of the prototype.
@@
The visualization location change VLC, previously introduced as the prototypical part location change by Sacha et al. [63], assesses the change in the location of the bounding box following a perturbation by calculating the intersection over union.
@@
To evaluate the similarity change, we use the prototype similarity change PSC, formerly introduced as prototype activation change by Sacha et al. [63]. This metric measures the relative change in the similarity score after the perturbation is applied.
$\mathit{PSC}=\frac{|(s_{p_{i}}(x)-s_{p_{i}}\left(\overline{x}_{i}\right)|}{s_{p_{i}}(x)}$ (7)
@@
We also assess location and activation changes in the feature space using the similarity map. To measure the location change, we introduce the following two metrics. The first metric is the prototype location change PLC, which measures the change in the maximum activation location between two similarity maps. We will use the Manhattan distance for this measure.
@@
The other metric is the prototype activation location change PALC. This metric measures the location change of the hole activation pattern in the similarity map, analogous to the VLC metric. We apply a min-max normalization and use a threshold of $0.5$ to obtain a binary activation pattern, expressed by the $\operatorname{bin}()$ operation.
$\mathit{PALC}=1-\frac{\left|\operatorname{bin}(a_{p_{i}}(x))\cap\operatorname{bin}(a_{p_{i}}\left(\overline{x}_{i}\right))\right|}{\left|\operatorname{bin}(a_{p_{i}}(x))\cup\operatorname{bin}(a_{p_{i}}\left(\overline{x}_{i}\right))\right|}$ (10)
To evaluate the activation change in similarity maps, we introduce the prototype activation change PAC, which operates similar to the VAC metric. Let $m\times n$ represent the height and width of the feature map.
$\mathit{PAC}=1-\frac{\sum_{j=1}^{m\times n}\min(a_{p_{i}}(x)_{j},a_{p_{i}}(\overline{x}_{i})_{j})}{\sum_{j=1}^{m\times n}\max(a_{p_{i}}(x)_{j},a_{p_{i}}(\overline{x}_{i})_{j})}$ (11)
@@ 5.2 Continuity Evaluation
The continuity evaluation assesses the stability of prototypes under different augmentations. We focus on the study by Hoffmann et al. [51], which assesses prototype stability under noise and compression artifacts, and the studies by Rymarczyk et al. [34] and Nauta et al. [57], which examine the influence of photometric augmentations. We train our model using geometric and photometric augmentations to promote continuity during training, as recommended by Nauta et al. [19]. In the evaluation, we perturb the input image $x$ to $\overline{x}$ with photometric and noise augmentations. The individual augmentations are applied all at once with the following settings: brightness $12.5\%$, contrast $12.5\%$, saturation $12.5\%$, hue $0.05$, noise $5\%$, JPEG quality $90\%$, and blur using a $3\times 3$ kernel. A visual presentation of the process is illustrated in Figure 1.
We utilize the metrics from the output-completeness evaluation to evaluate the changes in location and activation within the major components of the models (Figure 1). The VLC and VAC scores are excluded, as they focus on assessing the visualization method instead of a prototype. In addition, we use the prototype rank change PRC from Sacha et al. [63]. This metric measures the prototype’s position change in the similarity score ranking of all prototypes.
@@
The other metric is the classification rank change CRC that measures the change in the rank of the predicted class. Let $r_{c}(x)$ be the rank of the predicted class $c$ for the image $x$.
@@ 5.3 Contrastivity
The contrastivity evaluation is used to assess the differences between prototypes. To evaluate the contrastivity of prototypes, we refer to the study by Wang et al. [64], which evaluated the contrastivity of prototypes in the feature space. We follow the recommendations of Nauta et al. [19] to also assess location differences.
Wang et al. [64] measured the average cosine distance between prototypes $\mathit{APD_{inter}}$ and feature vectors $\mathit{AFD_{inter}}$ of different classes. These metrics are designed for networks with class-specific prototypes, so each class has a subset of prototypes $P_{k}\subset P$, where $P$ is the set of all prototypes and $k$ is the class index. To ensure compatibility with models that do not assign prototypes to classes, we use the ground truth class of the image $x$ to create the subset $P_{x}$, consisting of the top-5 most activated prototypes in the image. The subset $P_{k}$ is then the combination of all subsets $P_{x}$ in the test set $X_{test}$. To create the corresponding feature vector sets $F_{k}$, we use the nearest feature vectors of the prototypes in the subset $P_{x}$.
We use the average inter-class distance metrics from Wang et al. [64] and also employ the average intra-class distance case $\mathit{APD_{intra}}$ and $\mathit{AFD_{intra}}$ respectively. Following the recommendations from Nauta et al. [19], we use the PLC and PALC metrics to assess the contrastivity of prototype locations in the feature map and introduce a new metric to evaluate the activation discriminativeness of prototypes, achieving a comprehensive model coverage (see Figure 1).
The original study from Wang et al. [64] does not state the specific formula of the metrics, so the following calculations are our interpretation of the metrics. The average inter-class prototype distance $\mathit{APD_{inter}}$ from Wang et al.  [64] is the average cosine distance between prototypes of different classes.
@@
We calculate the average inter-class feature distance $AFD_{inter}$ from Wang et al. [64] in the same way using the feature vector set $F$ instead of the prototype set $P$. To evaluate the contrastivity between prototypes from the same class, we introduce the average intra-class prototype distance $\mathit{APD_{intra}}$. The metric measures the average cosine distance of a prototype to all other prototypes of the same class.
@@
We also extend this metric to the feature vector case introducing the average intra-class feature distance $\mathit{AFD_{intra}}$, by switching the prototype set $P$ with the feature vector set $F$. To measure the class discriminativeness of a prototype, we compute the Shannon entropy of the similarity scores. Let $S_{p_{i}}$ be the set of similarity scores of a prototype over all test images in $X_{test}$. We use a max normalization to normalize the similarity scores and then calculate the histogram $hist(S_{p_{i}})$ of the similarity scores with $U=10$ bins.
@@ 5.4 Covariate Complexity
The covariate complexity evaluation assesses the complexity of a prototype regarding the interpretability. To evaluate the covariate complexity of prototypes, we focus on the study by Wang et al. [64], which evaluated the overlap of prototype visualizations with object masks, and the study by Huang et al.  [53], which evaluated how consistently object parts are represented by a prototype.
The following metrics from Wang et al. [64] are again our interpretation, as the study does not provide specific formulas. To determine the activated region of a prototype in the input image, we again use the 95th percentile approximation on the prototype visualization for these metrics.
To assess the overlap between prototypes and objects, Wang et al. [64] introduces the content heatmap metric. We renamed this metric to $\mathit{Object\ Overlap}$ to be consistent with our naming scheme. The metric calculates the intersection of the prototype visualization and the object mask. Let $m(x)$ be the object mask of image $x$.
@@
We also adapted the idea of the outside-inside relevance ratio measure from Wang et al. [64], introducing the inside-outside relevance difference IORD as the signed difference between mean inside and outside activation. The adoption of the original metric was made to increase the stability of the measurements. Each saliency map is normalized before the calculation to ensure that IORD scores are consistent between different saliency maps.
@@
To measure the alignment of prototypes with specific object parts, we adapted the consistency score from Huang et al. [53]. The coverage of an object part is calculated by computing a histogram of the object parts covered by a prototype over the test set $X_{test}$. This is done by computing the bounding box and adding all object parts within the bounding box to the histogram of that prototype. After processing every image in the test set, we normalize the histograms. This results in a vector with object part percentages, denoted as $l_{p_{i}}$. The consistency of the prototype is the average coverage measure over the vector $l_{p_{i}}$.
@@ 5.5 Compactness
The compactness evaluation assesses general characteristics of the model. This evaluation will focus on the study by Nauta et al. [21], which evaluated global size, local size, and the sparsity of the classification layer. Additionally, we will include the negative-positive reasoning ratio metric in our compactness evaluation.
The global explanation size from Nauta et al. [21] counts all prototypes with at least one non-zero weight in the classification layer. This definition is not directly applicable to the ProtoPool model. Therefore, we will count non-zero weights in the prototype presence matrix from ProtoPool. The local explanation size, as described by Nauta et al. [21], can be defined as follows. Let $\hat{s}_{p_{i}}(x)$ be the normalized similarity score of a prototype $p_{i}$ for an image $x$, with $s(x)$ being the similarity score vector of all prototypes for the image $x$.
@@
The classification Sparsity from Nauta et al. [21] is the ratio of zero weights in the classification layer. Let $W$ represent the weights in the classification layer, and $\epsilon=0.001$ be the threshold for considering a weight as non-zero to improve the numerical stability.
@@
To evaluate the positive reasoning property of the model, we introduce the negative-positive reasoning ratio NPR, that calculates the ratio between the number of positive and negative weights in the classification layer. We will again use a threshold of $\epsilon=0.001$ for considering a weight as positive or negative.
@@ 6 Experimental Setup
In the original studies, datasets were split into 50% training and 50% test data. We will follow a different method based on the study by Raschka et al. [65], which argues that using the same data subset for model selection and final evaluation can result in overly optimistic outcomes. Following these guidelines, we first divide the datasets into 70% training and 30% test sets. We then apply 4-fold stratified cross-validation to split the training set into training and validation subsets, maintaining a consistent class distribution across all sets. This results in four different sets of training and validation pairs, with a 52.5% training and 17.5% validation split.
We selected the ResNet-50 architecture with pretrained ImageNet weights as a backbone for all models. The feature maps’ spatial dimensions are $7\times 7$. In the original study from PIPNet the dimension is modified to $28\times 28$. However, it is not described as a core architecture design, so we argue that the use of a uniform dimension setting makes the comparison fairer. For the dimension of explicit prototype vectors in the ProtoPNet and ProtoPool models, we chose $1\times 1\times 128$. The last layer of ProtoPNet is initialized with a 1 indicating a prototype belongs to a class and 0 otherwise instead of 0.5-0.5, because the sparsity objective described by Chen et al. [16] could otherwise not be achieved in our experiments. PIPNet uses 2048 prototypes for all datasets. ProtoPNet uses 2000, 1960, 50 and 490 for CUB200, Cars196, NICO and AWA2 respectively. ProtoPool uses 205, 201, 50, 168 prototypes for the datasets, respectively.
We used an online augmentation setting with the Albumentations library [66], that supports augmenting bounding boxes, segmentation masks, and key-points along with images, which is crucial for certain metrics in our evaluation. All networks were trained with geometric and photometric augmentations. The original studies from ProtoPNet and ProtoPool only used geometric augmentations, but due to our consistency evaluation, we choose to add additional photometric augmentations to promote better results. In the original study, the PIPNet model used photometric augmentation for the contrastive learning approach, which is also included in our experiments. The experiments on the CUB200 and Cars196 were done on cropped images using the bounding box information.
@@
Due to the differences between our training setup and the original studies, we adjusted the learning rates, training epochs, and schedulers for each model and dataset. The new parameters were chosen based on hyperparameter tuning with 50 trails for each dataset. To reduce complexity and the computational load, we used a fixed number of 100 joint epochs to tune all models. The individual loss weights were kept the same as the original paper suggested over all datasets. All networks were re-implemented based on the original code for our experiments and are also part of our open-source library. We point out that our training setup deviates from the optimal learning procedure introduced in the original studies, and we expect a reduction in performance. Therefore, our focus is to evaluate if interpretable properties discussed in the original studies remain in suboptimal learning conditions.

### v3

@@ 4.1 Prototype Visualization
The post-hoc visualization method is not bound to the individual models, allowing us to select the most suitable method for the networks. The original implementation of ProtoPNet employs a simple model-agnostic upscaling method, which was also used by ProtoPool and PIPNet. This method upscales the similarity map of a prototype to the size of the input using a cubic interpolation to visualize the similarity pattern of the prototype within the image. A key disadvantage is that it relies solely on the similarity map and assumes the preservation of spatial information in the feature extraction process. There are other model-specific methods that can better visualize the feature extraction process of the model. Based on these arguments and the findings from Xu-Darme et al.Xu-Darme et al. [2023b], which compared various visualization methods, we selected the Prototypical Relevance Propagation (PRP) method from Gautam et al.Gautam et al. [2023] for visualizing the prototypes (see Figure 1).
@@ 5 Evaluation Metrics
In this section, we discuss the techniques and metrics used to evaluate the part-based prototype models. Following the Co-12 properties from Nauta et al. Nauta and Seifert [2023] as a guideline, we choose metrics to evaluate the output-completeness, continuity, contrastivity, covariate complexity, and compactness.
@@
We use a total of 22 metrics, including 13 metrics that are novel to this domain. Figure 1 provides an overview of which metrics address which model components (A-G). Additional categories indicate whether a metric measures location, activation, or semantic properties. Some metrics are used in multiple evaluation techniques, indicated by a subscript. An overview of which metric is used in which technique, with a brief description of how the metric should be interpreted in each technique, can be found in Table 1.
@@ 5.1 Output-Completeness Evaluation
The output-completeness evaluation focuses on the visualization method’s ability to highlight all relevant parts in the input image. To evaluate the output completeness of the PRP method, we consider the study by Sacha et al. Sacha et al. [2023b], which evaluates the spatial misalignment of visualization methods. The introduced metrics measure the location change in the visualization VLC and the prototype activation change PSC when a perturbation is applied to the input image. We extend this list by including the activation change of the visualization VAC, the location change of the maximum activation in the feature map PLC, as well as the general location and activation change of the similarity pattern, PALC and PAC, respectively, thereby achieving comprehensive coverage of the model components, see Figure 1. Some metrics from Sacha et al. Sacha et al. [2023b] have been renamed for consistency in terminology. Furthermore, some metrics are also used in continuity and contrastivity evaluation with different input variables. Here, the according formulas are only presented once.
Consider an image $x$ and the perturbed counterpart $\overline{x}_{i}$, created based on the visualization of the prototype $p_{i}\in P$. Let $v_{p_{i}}(x)$ be the saliency map of prototype $p_{i}$. Each saliency map is pre-processed and only contains pixels with a relevance value above the 95th percentile. Let $b_{p_{i}}(x)$ be the bounding box around this activated region of the saliency map. For the output-completeness evaluation, we perturb an image $x$ to $\overline{x}_{i}$ by adding a Gaussian noise mask with a standard deviation of $0.05$ to the pixels around the bounding box; see Figure 5 for a process visualization. Further, let $a_{p_{i}}(x)$ be the similarity map and $s_{p_{i}}(x)$ the similarity score of the prototype.
@@
The visualization location change VLC, previously introduced as the prototypical part location change by Sacha et al. Sacha et al. [2023b], assesses the change in the location of the bounding box following a perturbation by calculating the intersection over union.
@@
To evaluate the similarity change, we use the prototype similarity change PSC, formerly introduced as prototype activation change by Sacha et al. Sacha et al. [2023b]. This metric computes the difference between two similarity values and is used to measure the relative change in the similarity score after the perturbation is applied.
$\mathit{PSC}_{out}=\frac{|(s_{p_{i}}(x)-s_{p_{i}}\left(\overline{x}_{i}\right)|}{s_{p_{i}}(x)}$ (7)
@@
We also assess location and activation changes in the feature space using the similarity map. In order to measure the location change, we introduce two metrics: First, the prototype location change PLC, measuring the distance between two maximum activation locations using the Manhattan distance. The output-completeness evaluation uses the similarity maps of a prototype from the original and perturbed image version.
@@
The second metric is the prototype activation location change PALC. This metric measures the location difference of the hole activation pattern between two similarity maps, analogous to the VLC metric. We apply a min-max normalization and use a threshold of $0.5$ to obtain a binary activation pattern, expressed by the $\operatorname{bin}()$ operation. Again, the output-completeness evaluation uses the original and perturbed images to generate the two similarity maps.
$\mathit{PALC}_{out}=1-\frac{\left|\operatorname{bin}(a_{p_{i}}(x))\cap\operatorname{bin}(a_{p_{i}}\left(\overline{x}_{i}\right))\right|}{\left|\operatorname{bin}(a_{p_{i}}(x))\cup\operatorname{bin}(a_{p_{i}}\left(\overline{x}_{i}\right))\right|}$ (10)
In order to evaluate the activation change in similarity maps, we introduce the prototype activation change PAC, which operates similar to the VAC metric and measures the activation difference between two similarity maps. Let $m\times n$ represent the height and width of the feature map, then
$\mathit{PAC}_{out}=1-\frac{\sum_{j=1}^{m\times n}\min(a_{p_{i}}(x)_{j},a_{p_{i}}(\overline{x}_{i})_{j})}{\sum_{j=1}^{m\times n}\max(a_{p_{i}}(x)_{j},a_{p_{i}}(\overline{x}_{i})_{j})}$ (11)
@@ 5.2 Continuity Evaluation
The continuity evaluation assesses the stability of prototypes under different augmentations. We focus on the study by Hoffmann et al. Hoffmann et al. [2021], which assesses prototype stability under noise and compression artifacts, and the studies by Rymarczyk et al. Rymarczyk et al. [2021] and Nauta et al. Nauta et al. [2021b], which examine the influence of photometric augmentations. We train our model using geometric and photometric augmentations to promote continuity during training, as recommended by Nauta et al. Nauta and Seifert [2023]. In the evaluation, we perturb the input image $x$ to $\overline{x}$ with photometric and noise augmentations. The individual augmentations are applied all at once with the following settings: brightness $12.5\%$, contrast $12.5\%$, saturation $12.5\%$, hue $0.05$, noise $5\%$, JPEG quality $90\%$, and blur using a $3\times 3$ kernel. A visual presentation of the process is illustrated in Figure 1.
We utilize the metrics from the output-completeness evaluation to evaluate the changes in location and activation within the major components of the models (Figure 1). This is done by switching the prototype-specific perturbation $\overline{x}_{i}$ with the general perturbation $\overline{x}$ based on the specified augmentation techniques. The metrics adopted are PLC, PSC, PALC, PRC, and PAC. The VLC and VAC scores are excluded, as they focus on assessing the visualization method instead of a prototype. In addition, we use the prototype rank change PRC from Sacha et al. Sacha et al. [2023b]. This metric measures the prototype’s position change in the similarity score ranking of all prototypes.
@@
The other metric is the classification rank change CRC that measures the change in the rank of the predicted class. Let $r_{c}(x)$ be the rank of the predicted class $c$ for the image $x$:
@@ 5.3 Contrastivity
The contrastivity evaluation is used to assess the differences between prototypes. To evaluate the contrastivity of prototypes, we refer to the study by Wang et al. Wang et al. [2023], which evaluated the contrastivity of prototypes in the feature space. We follow the recommendations of Nauta et al. Nauta and Seifert [2023] to also assess location differences.
Wang et al. Wang et al. [2023] measured the average cosine distance between prototypes $\mathit{APD_{inter}}$ and feature vectors $\mathit{AFD_{inter}}$ of different classes. These metrics are designed for networks with class-specific prototypes, so each class has a subset of prototypes $P_{k}\subset P$, where $P$ is the set of all prototypes and $k$ is the class index. To ensure compatibility with models that do not assign prototypes to classes, we use the ground truth class of the image $x$ to create the subset $P_{x}$, consisting of the top-5 most activated prototypes in the image. The subset $P_{k}$ is then the combination of all subsets $P_{x}$ in the test set $X_{test}$. To create the corresponding feature vector sets $F_{k}$, we use the nearest feature vectors of the prototypes in the subset $P_{x}$.
We use the average inter-class distance metrics from Wang et al. Wang et al. [2023] and also employ the average intra-class distance case $\mathit{APD_{intra}}$ and $\mathit{AFD_{intra}}$ respectively. Following the recommendations from Nauta et al. Nauta and Seifert [2023], we use the PLC and PALC metrics to assess the contrastivity of prototype locations in the feature map and introduce a new metric to evaluate the activation discriminativeness of prototypes, achieving a comprehensive model coverage (see Figure 1). To use the PLC and the PALC described in the output-completeness evaluation in this context, we change the variable $a_{p_{i}}(\overline{x}_{i})$ to $a_{p_{j}\in P_{x}/p_{i}}(x)$ in order to compare the similarity maps between activated prototypes in image $x$
The original study from Wang et al. Wang et al. [2023] does not state the specific formula of the metrics, so the following calculations are our interpretation. The average inter-class prototype distance $\mathit{APD_{inter}}$ from Wang et al.  Wang et al. [2023] is the average cosine distance between prototypes of different classes.
@@
We calculate the average inter-class feature distance $\mathit{AFD_{inter}}$ from Wang et al. Wang et al. [2023] in the same way using the feature vector set $F$ instead of the prototype set $P$. To evaluate the contrastivity between prototypes from the same class, we introduce the average intra-class prototype distance $\mathit{APD_{intra}}$. The metric measures the average cosine distance of a prototype to all other prototypes of the same class.
@@
We also extend this metric to the feature vector case via introducing the average intra-class feature distance $\mathit{AFD_{intra}}$, by switching the prototype set $P$ with the feature vector set $F$. In order to measure the class discriminativeness of a prototype, we compute the Shannon entropy of the similarity scores. Let $S_{p_{i}}$ be the set of similarity scores of a prototype over all test images in $X_{test}$. We use a max normalization to normalize the similarity scores and then calculate the histogram $hist(S_{p_{i}})$ of the similarity scores with $U=10$ bins.
@@ 5.4 Covariate Complexity
The covariate complexity evaluation assesses the complexity of a prototype regarding the interpretability. To evaluate the covariate complexity of prototypes, we focus on the study by Wang et al. Wang et al. [2023], which evaluated the overlap of prototype visualizations with object masks, and the study by Huang et al.  Huang et al. [2023], which evaluated how consistently object parts are represented by a prototype.
The following metrics from Wang et al. Wang et al. [2023] are again our interpretation, as the study does not provide specific formulas. To determine the activated region of a prototype in the input image, we again use the 95th percentile approximation on the prototype visualization for these metrics.
To assess the overlap between prototypes and objects, Wang et al. Wang et al. [2023] introduces the content heatmap metric. We renamed this metric to $\mathit{Object\ Overlap}$ to be consistent with our naming scheme. The metric calculates the intersection of the prototype visualization and the object mask. Let $m(x)$ be the object mask of image $x$.
@@
We also adapted the idea of the outside-inside relevance ratio measure from Wang et al. Wang et al. [2023], introducing the inside-outside relevance difference IORD as the signed difference between mean inside and outside activation. The adoption of the original metric was made to increase the stability of the measurements. Each saliency map is normalized before the calculation to ensure that IORD scores are consistent between different saliency maps.
@@
In order to measure the alignment of prototypes with specific object parts, we adapted the consistency score from Huang et al. Huang et al. [2023]. The coverage of an object part is calculated by computing a histogram of the object parts covered by a prototype over the test set $X_{test}$. This is done by computing the bounding box and adding all object parts within the bounding box to the histogram of that prototype. After processing every image in the test set, we normalize the histograms. This results in a vector with object part percentages, denoted as $l_{p_{i}}$. The consistency of the prototype is the average coverage measure over the vector $l_{p_{i}}$.
@@ 5.5 Compactness
The compactness evaluation assesses general characteristics of the model. This evaluation will focus on the study by Nauta et al. Nauta et al. [2023a], which evaluated global size, local size, and the sparsity of the classification layer. Additionally, we will include the negative-positive reasoning ratio metric in our compactness evaluation.
The global explanation size from Nauta et al. Nauta et al. [2023a] counts all prototypes with at least one non-zero weight in the classification layer. This definition is not directly applicable to the ProtoPool model. Therefore, we will count non-zero weights in the prototype presence matrix from ProtoPool. The local explanation size, as described by Nauta et al. Nauta et al. [2023a], can be defined as follows. Let $\hat{s}_{p_{i}}(x)$ be the normalized similarity score of a prototype $p_{i}$ for an image $x$, with $s(x)$ being the similarity score vector of all prototypes for the image $x$.
@@
The classification Sparsity from Nauta et al. Nauta et al. [2023a] is the ratio of zero weights in the classification layer. Let $W$ represent the weights in the classification layer, and $\epsilon=0.001$ be the threshold for considering a weight as non-zero to improve the numerical stability.
@@
In order to evaluate the positive reasoning property of the model, we introduce the negative-positive reasoning ratio NPR, that calculates the ratio between the number of positive and negative weights in the classification layer. We will again use a threshold of $\epsilon=0.001$ for considering a weight as positive or negative.
@@ 6 Experimental Setup
In the original studies, datasets were split into 50% training and 50% test data. We will follow a different method based on the study by Raschka et al. Raschka [2018], which argues that using the same data subset for model selection and final evaluation can result in overly optimistic outcomes. Following these guidelines, we first divide the datasets into 70% training and 30% test sets. We then apply 4-fold stratified cross-validation to split the training set into training and validation subsets, maintaining a consistent class distribution across all sets. This results in four different sets of training and validation pairs, with a 52.5% training and 17.5% validation split.
We selected the ResNet-50 architecture with pretrained ImageNet weights as a backbone for all models. The feature maps’ spatial dimensions are $7\times 7$. In the original study from PIPNet the dimension is modified to $28\times 28$. However, it is not described as a core architecture design, so we argue that the use of a uniform dimension setting makes the comparison fairer. For the dimension of explicit prototype vectors in the ProtoPNet and ProtoPool models, we chose $1\times 1\times 128$. The last layer of ProtoPNet is initialized with a 1 indicating a prototype belongs to a class and 0 otherwise instead of 0.5-0.5, because the sparsity objective described by Chen et al. Chen et al. [2019] could otherwise not be achieved in our experiments. PIPNet uses 2048 prototypes for all datasets. ProtoPNet uses 2000, 1960, 50 and 490 for CUB200, Cars196, NICO and AWA2 respectively. ProtoPool uses 205, 201, 50, 168 prototypes for the datasets, respectively.
We used an online augmentation setting with the Albumentations library Buslaev et al. [2020], that supports augmenting bounding boxes, segmentation masks, and key-points along with images, which is crucial for certain metrics in our evaluation. All networks were trained with geometric and photometric augmentations. The original studies from ProtoPNet and ProtoPool only used geometric augmentations, but due to our consistency evaluation, we choose to add additional photometric augmentations to promote better results. In the original study, the PIPNet model used photometric augmentation for the contrastive learning approach, which is also included in our experiments. The experiments on the CUB200 and Cars196 were done on cropped images using the bounding box information.
@@
Due to the differences between our training setup and the original studies, we adjusted the respective learning rates, training epochs, and schedulers for each model and dataset. The new parameters were chosen based on hyperparameter tuning with 50 trails for each dataset. To reduce complexity and the computational load, we used a fixed number of 100 joint epochs to tune all models. The individual loss weights were kept the same as the original paper suggested over all datasets. All networks were re-implemented based on the original code for our experiments and are also part of our open-source library. We point out that our training setup deviates from the optimal learning procedure introduced in the original studies, and we expect a reduction in performance. Therefore, our focus is to evaluate if interpretable properties discussed in the original studies remain in suboptimal learning conditions.